OpenStudy (pokiedokie):

|3x-2|+4 < (less than or equal to) 7 1/2|x+2|>6 can you please answer these with a guide/steps through it? i'm doing a study guide and need a little help

OpenStudy (lynfran):

wat do u have to do here?

OpenStudy (lynfran):

and r those separate equality

OpenStudy (lynfran):

?

OpenStudy (pokiedokie):

solve the inequality, and then i can graph the solution on my own

OpenStudy (pokiedokie):

yes

OpenStudy (lynfran):

ok

OpenStudy (lynfran):

\[|3x-2|+4\le7\]\[|3x-2|+4-4\le 7-4\]\[|3x-2|\le7-4\]\[|3x-2|\le3\]\[(3X-2)\le \pm 3\]

OpenStudy (lynfran):

\[3x-2 \le 3\]\[3x-2+2 \le 3+2\]\[3x \le 5\]\[x \le 5/3\]

OpenStudy (lynfran):

and \[3x-2 \le -3\]\[3x-2+2 \le -3+2\]\[3x \le -1\]\[x \le -1/3\]

OpenStudy (pokiedokie):

thank you so much!! this helps a lot

OpenStudy (lynfran):

ok for the 2nd inequality you can start by dividing both sides by 1/2...since 1/2 is being multiply to the absolute sign...

OpenStudy (lynfran):

u will have |x+2|>12 and the steps are like the first inequality

OpenStudy (pokiedokie):

ahh okay, i see. thank you

OpenStudy (lynfran):

welcome